OpenStudy (anonymous):

g(x)=ln(x^2)/(x) how does squaring "x" on the top and then having "x" as the denominator affect the different portions of the graph?

OpenStudy (anonymous):

\[ln(x^2)=2ln(x)\]

OpenStudy (anonymous):

so your function is \[\frac{2\ln(x)}{x}\]

OpenStudy (anonymous):

i don't know that this answers your question or not

OpenStudy (anonymous):

no i need to know see how on the top ln(x) is multiplied by 2 and then it has a denominator of x, how just like before in my last question do those 2 things affect the different portions of its graph